RK Az

The parallel circuit above contains 2 resistors, each with a resistance R. If a third resistor (Rs) is connected in parallel with the other two resistors, what will happen to current measured by the ammeters? Ignore the resistance of the wires.

  1. A1 will increase and A2 will stay the same

  2. A1 and A2 will stay the same

  3. A1 will increase and A2 will decrease

  4. A1 and A2 will increase
Answer #1

option A1 and A2 will increase is correct.

Reason: when one resister is there the I =v/r will give the current

when two resistors are connected in parallel the I =V/r =v/(R/2)= 2(v/r) current value increases two times the circuit contains one resistor.

similarly in three resistor containing circuit the current value will increase than the circuit containing two resistors in parallel. so the current in two equivalent resistor remains same due to same resistor and current through the Rs depends on the value of Rs.